OpenStudy (anonymous):

Which of the following is a solution of 4x2 = - 9x - 4?

OpenStudy (anonymous):

First things first, you may want to bring the terms on the right over.

OpenStudy (anonymous):

done i plug it in and everything but i get. 

OpenStudy (anonymous):

@Indivicivet

OpenStudy (anonymous):

\[b^2-4ac = 9^2-4^3 = 81-64 = 17\]Looks good to me!

OpenStudy (anonymous):

but its not on the answers @Indivicivet

OpenStudy (anonymous):

What are the answers? Did you make a mistake transcribing somewhere or did we both manage to mess up? ^^

OpenStudy (anonymous):

Which of the following is a solution of 4x2 = - 9x - 4?

OpenStudy (anonymous):

@Indivicivet

OpenStudy (anonymous):

The last one matches up and the "x-" looks like a formatting error to me. ^^;

OpenStudy (anonymous):

i hope so thanks!! Can i show you another one i'm stuck in?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

the one in the bottom is my answer

OpenStudy (anonymous):

What kind of baseball player hits a ball straight up?! Anyway... \[0 = -16t^2 + 64t + 4 = 4t^2 - 16t - 1\] Does that help?

OpenStudy (anonymous):

haha yeah. I'm gonna try it now thanks!

OpenStudy (anonymous):

didn't really help :/ @Indivicivet

OpenStudy (anonymous):

Using the quadratic formula, I get \[2 \pm 4\sqrt{17}\]in my head, using the answers for minor guidance. You seemed fine on the first part; do you know what's tripping you up here?